# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ 4.1 Approximating and solving the OT problem for GRFs
OT, permutation densities and bipartite matchings. The OT problem in Eq. 13 is analytically intractable. To make progress, we must make some approximations. We report them in full in App. C.3; Fig. 6 provides a visual overview. We limit the main text to the most important points.
Our first problem is intractability of the cost function in Eq. 13. We can mitigate this by moving $\mathbb{E}_{\textrm{dirs}}$, the expectation over walker directions, inside the square brackets. This clearly modifies the objective, ignoring higher-order correlations between features induced by walker directions (but not lengths). However, it yields a cost function we can estimate efficiently, if not compute exactly.
The second challenge is optimising over the class of couplings $\Lambda_{2}(\eta_{\textrm{G}})$, joint distributions over two discrete random variables with geometrically distributed marginals. As in Sec. 3.2, a sensible numerical approach is to optimise amongst a tractable subclass of $\Lambda_{2}(\eta_{\textrm{G}})$. Copulas are unsuitable because in this discrete setting we cannot compute gradients to optimise the covariance matrix. Instead, we consider the family of measures $\pi^{(\sigma)}$ on $[0,1]^{2}$ described by the permutation densities
$p_{\sigma}(x,y)\coloneqq n1_{\sigma(\lceil nx\rceil)=\lceil ny\rceil},$ (14)
with $\sigma$ a permutation of order $n$ (that is, a bijection $\sigma:[\![n]\!]\to[\![n]\!]$). The unit square is split into a $n\times n$ grid where each row and column has a single ‘tile’ of probability density $n$ and is 00 otherwise.
Note that both marginal distributions of $\pi^{(\sigma)}$ are manifestly uniform on $[0,1]$. Moreover, the marginals can be transformed to an arbitrary probability distribution $\eta$ by pushing forward with the inverse CDF, $F_{\eta}^{-1}(\cdot)\coloneqq\inf\{x\in\mathbb{R}:F_{\eta}(x)\geq\cdot\}$. Transforming both coordinates in this way yields a joint measure $\mu^{(\sigma)}\in\Lambda_{2}(\eta)$ which will give an unbiased estimator. We will refer to these as $\sigma$-couplings, and remark that $n!$ such couplings exist for a given permutation order $n$.
The permutations $\sigma\in S_{n}$ can be interpreted as matchings between the $n$ quantiles of the geometric distributions over the lengths of a pair of walkers. With the right choice of $\sigma$, they can ensure that e.g. if one of the walk lengths is short then the other tends to be long, diversifying the ensemble. In App. C.1 we formalise this by writing the optimisation of $\sigma$ as a bipartite matching problem, with a cost matrix estimated by sampling walks. Unfortunately, the matching problem has quadratic weights (albeit with some extra symmetric structure). It is challenging; we discuss progress towards a fast, exact algorithm in App. C.3.1. More pragmatically, with another simplification we can obtain a related linear-weights bipartite matching problem. This can be solved efficiently in time $\mathcal{O}(n^{3})$ using linear programming techniques, e.g. the Hungarian algorithm (Kuhn, 1955), thereby optimising the coupling $\mu^{(\sigma)}$. Taking the optimal permutation $\sigma$, we define $\sigma$-coupled GRFs as follows.
@@
We refer to GRFs as $\sigma$-coupled if they are constructed using pairs of random walks with lengths drawn from the coupling $\mu^{(\sigma)}$, with the optimal permutation $\sigma$ obtained by solving a matching problem between the quantiles of the distributions over walk lengths (see App. C.3 for details).
Besides being easy to optimise and sample from, there are also more rigorous OT motivations for the choice of $\sigma$-couplings $\mu^{(\sigma)}$. They relate to the asymptotic behaviour of $\mu^{(\sigma)}$ as the permutation order $n\to\infty$ and the stability of OT plans (Villani, 2021). We defer this technical point to App. D.
As in Sec. 3.2, an interesting question is whether one could couple the lengths of $m>2$ walkers. This is challenging and has received little attention in the literature. One possibility would be to combine $m-1$ permutations, finding a minimum-weights $m$-partite matching with all subgraphs constrained to be complete $K_{m}$. Another approach would be to approximately solve this multi-marginal OT problem using the Sinkhorn-Knopp algorithm (Sinkhorn and Knopp, 1967; Cuturi, 2013).
Broader applicability. As a final remark, the utility of our OT-approach to coupling walker lengths extends to graph-based estimators beyond GRFs. For example, it can be used to improve estimates of the PageRank vector, a popular measure of the relative importance of the nodes $\mathcal{N}$ of a graph proposed by Page et al. (1998) to rank websites in search engine results. We discuss this in App. E.
@@ 4.2 Experiments with $\sigma$-coupled GRFs
In this section, we empirically evaluate $\sigma$-coupled GRFs for variance reduction of graph node kernel estimates. We begin by testing our method on a variety of real-world graphs, then show that lower variance unlocks better approximate inference with scalable graph-based Gaussian processes (GPs).
Figure 1: Kernel estimator error vs termination probability. Insets show permutations for $p\in\{0.1,0.2,0.3,0.4,0.5\}$.
@@
Gram matrix approximation. GRFs take the termination probability $p$ as a hyperparameter, determining the rate of decay of the geometric distribution over walk length. A smaller value of $p$ samples longer walks and gives more accurate kernel estimates, but takes longer to run. The optimal coupling changes depending on $p$. We will consider the values $p\in\{0.1,0.2,0.3,0.4,0.5\}$, finding the optimal permutation $\sigma$ in each case.
To find the optimal $\sigma$, we solve the matching problem (App. C.3) for a random Erdős-Rényi graph with $N=100$ nodes, taking a permutation order $n=30$ and choosing the $2$-regularised Laplacian kernel as our target. We then use the Hungarian algorithm, averaging the cost matrix over every possible node pair $(v_{i},v_{j})\in\mathcal{N}^{2}$ to find a coupling that reduces the variance of all entries of the Gram matrix.
@@
Having computed couplings $\mu^{(\sigma)}$ for each value of $p$, we then test the corresponding $\sigma$-coupled GRFs on a variety of real-world graphs. Fig. 1 shows the results for cora ($N=2708$), with the rest left to App. F.1. We plot the relative Frobenius norm error of the Gram matrix approximation $\|\mathbf{K}-\widehat{\mathbf{K}}\|_{\textrm{F}}/\|\mathbf{K}\|_{\textrm{F}}$ with walkers that are i.i.d., antithetic (Reid et al., 2024c) or $\sigma$-coupled. For every value of $p$, $\sigma$-coupled GRFs give equally good or smaller kernel estimator errors. Our general-purpose OT approach often substantially outperforms antithetic termination: a bespoke algorithm designed specifically to improve GRFs. We include visualisations of the optimal permutations for different values of $p$ in the inset, verifying that the $\sigma$-coupling adapts to different GRF hyperparamaters.
@@
Novel application: $\sigma$-coupled GRFs for scalable graph-based GPs. We conclude this section by applying $\sigma$-coupled GRFs to scalable graph-based Gaussian processes (GPs), where we find that the improved estimation of the covariance function permits better approximate inference. Scalable GPs are a novel application of GRFs that may be of independent interest; we provide discussion in App. F.
@@
Let us consider the task of probabilistic graph interpolation. This aims to predict unknown values corresponding to particular graph nodes from a set of observed values, along with principled uncertainty estimates (Pfaff et al., 2020). We consider mesh graphs $\mathcal{G}$ where every node $v_{i}\in\mathcal{N}$ is associated with a normal vector $\bm{n}_{i}\in\mathbb{R}^{3}$ (Dawson-Haggerty, 2023). The objective is to predict the $z$-components of a masked set, $\{(\bm{n}_{i})_{z}\}_{i=1}^{N_{\textrm{test}}}$ with $N_{\textrm{test}}=\lfloor 0.05N\rfloor$. To predict the missing normal vectors, we use a graph-based GP with a heat kernel covariance function. We compute a sparse, unbiased approximation of this kernel using GRFs with $\{16,32,64\}$ walkers that are i.i.d., antithetic (Reid et al., 2024c) or $\sigma$-coupled. Details of GP hyperparameter optimisation are given in App. F.2.
Fig. 2 shows the results. For mesh graphs of different sizes (the largest as big as $8700$ nodes), we plot the relative Frobenius norm error of the Gram matrix approximation, the test root mean square error (RMSE), and the KL divergence to the true posterior. We see that the improvement in covariance function estimation unlocks better predictive accuracy and better uncertainty quantification.
@@
Figure 2: Graph-based GP regression results. $\sigma$-coupled GRFs give more accurate kernel estimates, better predictions and better uncertainty estimates. Estimates of standard deviations are shaded.
@@
Probabilistic interpolation of traffic data. To further demonstrate, we train a scalable graph-based GP on a traffic flow dataset of the highways of San Jose, California, curated by Borovitskiy et al. (2021) using data from Chen et al. (2001) and OpenStreetMap. The graph has $1016$ nodes, with the traffic speed available at $325$. With this sparse, noisy dataset, $\sigma$-coupled GRFs again give more accurate predictions and better uncertainty estimates. Full results are reported in App. F.3.
@@ 5 Discussion and outlook
OT provides a powerful paradigm for variance reduction with random features. It offers proof techniques and numerical algorithms for finding novel RF couplings on continuous and discrete input domains. Whilst the presence of variance reduction is unambiguous, the downstream benefits it provides are less clear. With GRFs for scalable GPs, variance reduction permits much better approximate inference (Sec. 4.2). With RFFs and RLFs, this is not the case (App. B.2). This surprising finding suggests that, though popular, variance reduction is not always the right goal.
Right framing, wrong cost function. The reason for this counterintuitive behaviour is that, even if a coupling guarantees lower variance pointwise kernel estimates $\{k(\bm{x}_{i},\bm{x}_{j})\}_{i,j=1}^{N}$, functions like the predictive mean and KL-divergence are highly nonlinear in these estimates. For example, they may involve inverting a Gram matrix. It is hard to predict how the bias and variance of these downstream quantities will depend on the distribution of the pointwise estimates; they are not guaranteed to improve. Couplings that reduce the variance of $k(\bm{x}_{i},\bm{x}_{j})$ and $k(\bm{x}_{i},\bm{x}_{k})$ often also modify their covariance, which may effect downstream estimators that combine them. This includes seemingly innocuous quantities like the KL divergence from the true to the approximate prior (App. B.2).
We posit that OT provides the right framing for the problem of coupling RFs, but sometimes pointwise kernel variance is the wrong cost function. This naive choice may not fully capture how the joint distribution over kernel estimates determines downstream performance. Coupling to optimise e.g. the spectral properties of $\widehat{\mathbf{K}}$ may prove better (Choromanski et al., 2018; Avron et al., 2017a). Fortunately, OT provides a suite of theoretical and numerical tools achieve this; one simply modifies the cost function in Eq. 2. We hope this research will spur further work in this exciting direction.
@@ 6 Contributions and acknowledgements
Relative contributions. IR conceptualised the project, proposed the coupling mechanisms in Defs 3.3 and 4.1, proved the major theoretical contributions, ran the GRF experiments and wrote the manuscript. SM designed and ran the RFF and RLF GP experiments (Sec. 3.3), helped shape the project’s direction and made core contributions to the text. KC acted as the senior lead, providing technical guidance and developing the algorithms for the matching problem in App. C.3.1. RET met frequently throughout the project, giving important advice and support. AW provided helpful discussion, supervision and feedback on the manuscript.
Acknowledgements and funding. IR acknowledges support from a Trinity College External Studentship. SM acknowledges funding from the Vice Chancellor’s and the George and Marie Vergottis scholarship of the Cambridge Trust, and the Qualcomm Innovation Fellowship. RET is supported by Google, Amazon, ARM, Improbable and an EPSRC grant EP/T005386/1. AW acknowledges support from a Turing AI fellowship under grant EP/V025279/1 and the Leverhulme Trust via CFI.

### v2

@@ 4.1 Approximating and solving the OT problem for GRFs
OT, permutation densities and bipartite matchings. The OT problem in Eq. 12 is analytically intractable. To make progress, we must make approximations. We report full details in App. C.3, limiting the main text to a high-level discussion of the main points.
First, to make the objective amenable to Monte Carlo approximation, we move $\mathbb{E}_{\textrm{dirs}}$ inside the square. This is because, unlike the expression in Eq. 12, $\mathbb{E}_{\textrm{dirs}}(\psi(\bm{\omega}_{1,2}^{(i,j)}))$ can be efficiently estimated by simulating random walks. Second, we must optimise amongst the class of couplings $\Lambda_{2}(\eta_{\textrm{G}})$, joint distributions of two discrete random variables with geometrically distributed marginals. As in Sec. 3.2, a sensible numerical approach is to limit oneself to a tractable subclass of $\Lambda_{2}(\eta_{\textrm{G}})$. Taking inspiration from numerical OT, consider the family of measures $\pi^{(\sigma)}$ on $[0,1]^{2}$ described by the permutation densities $p_{\sigma}(x,y)\coloneqq n1_{\sigma(\lceil nx\rceil)=\lceil ny\rceil}$, with $\sigma$ a permutation of order $n$ (that is, a bijection $\sigma:[\![n]\!]\to[\![n]\!]$). The unit square is split into a $n\times n$ grid where each row and column has a single ‘tile’ of probability density $n$ and is 00 otherwise (Fig. 2 left). Both marginal distributions of $\pi^{(\sigma)}$ are uniform on $[0,1]$ and are transformed to a probability distribution $\eta$ by pushing forward with the inverse CDF, $F_{\eta}^{-1}(\cdot)\coloneqq\inf\{x\in\mathbb{R}:F_{\eta}(x)\geq\cdot\}$. Transforming both coordinates in this way yields a joint measure $\mu^{(\sigma)}\in\Lambda_{2}(\eta)$ which will give an unbiased estimator. $n!$ such couplings exist for a given permutation order $n$; we aim to efficiently find the one with the lowest variance.
The permutations $\sigma\in S_{n}$ can be interpreted as matchings between the $n$ quantiles of the geometric distributions over the lengths of a pair of walkers (Fig. 2 centre). With the correct choice of $\sigma$, they can ensure that e.g. if one of the walk lengths is short then the other tends to be long, diversifying the ensemble (Fig. 2 right). Optimising $\sigma$, the approximate OT problem can be written
$\sigma^{*}=\arg\min_{\sigma\in S_{n}}\sum_{q_{1}\in[\![n]\!]}\sum_{q_{2}\in[\![n]\!]}\left[\left(\widehat{\psi}(q_{1}^{(i)})+\widehat{\psi}(\sigma(q_{1})^{(i)})\right)^{\top}\left(\widehat{\psi}(q_{2}^{(j)})+\widehat{\psi}(\sigma(q_{2})^{(j)})\right)\right]^{2}$ (13)
where $\widehat{\psi}(q^{(i)})\coloneqq\mathbb{E}_{u\sim\mathcal{U}((\frac{q-1}{n},\frac{q}{n}])}\left(\mathbb{E}_{\textrm{dirs}}{\psi}(F^{-1}_{\eta_{\textrm{G}}}(u)^{(i)})\right)$. $\mathcal{U}((a,b])$ is the uniform distribution on the interval $(a,b]$. Eq. 13 is a quadratic assignment problem (Finke et al., 1987; Burkard et al., 1998). This family is generally NP hard, but in our case the cost function has some convenient extra symmetric structure. In fact, the following is true.
Theorem 4.1 (Solving Eq. 13 exactly).
Given a set of vectors $\{\widehat{\psi}(q^{(i,j)})\}_{q=1}^{n}\subset\mathbb{R}^{N}$, Eq. 13 can be solved with high probability in time complexity independent of $N$. Moreover, in the special case where $i=j$, it can be solved in polynomial time in $n$ (under mild technical assumptions on the set).
Proof sketch. Details and definitions are in App. C.3.1. The time complexity can be made independent of the number of nodes $N$ by performing dimensionality reduction using the celebrated Johnson-Lindenstrauss transformation (Dasgupta et al., 2010), which preserves pairwise dot products with high probability. In the special case $i=j$, Eq. 13 can be rewritten as finding a permutation $\sigma$ that minimises the $L_{2}$ norm of some particular $N^{2}$-dimensional vector. In App. C.3.1 we provide a novel algorithm to achieve this efficiently by projecting onto a sequence of random Gaussian vectors, requiring only a mild geometrical condition called $\epsilon$-separation. See Lemma C.2 for full details. ∎
More pragmatically, one can set $q_{1}=q_{2}$ to simplify to a related linear assignment problem, which can be solved efficiently in time $\mathcal{O}(n^{3})$ using e.g. the Hungarian algorithm (Kuhn, 1955). We empirically investigate how this final approximation modifies the objective in App. C.3. Taking the optimal permutation and corresponding coupling $\mu^{(\sigma)}$, we define $\sigma$-coupled GRFs as follows.
@@
GRFs are $\sigma$-coupled if they are constructed using pairs of random walks with lengths drawn from the coupling $\mu^{(\sigma)}$, with the optimal permutation $\sigma$ obtained by solving a matching problem between the quantiles of the distributions over walk lengths (specifically, Eq. 56 in App. C.3).
Besides being easy to optimise and sample from, there are also more rigorous OT motivations for the choice of $\sigma$-couplings $\mu^{(\sigma)}$. They relate to the asymptotic behaviour of $\mu^{(\sigma)}$ as the permutation order $n\to\infty$ and the stability of OT plans (Villani, 2021). We defer this technical point to App. D. As in Sec. 3.2, another interesting question is whether one could couple the lengths of $m>2$ walkers. This is challenging and has received little attention in the literature. One possibility would be to combine $m-1$ permutations, finding a minimum-weights $m$-partite matching with all subgraphs constrained to be complete $K_{m}$. Another approach would be to approximately solve this multi-marginal OT problem using the Sinkhorn-Knopp algorithm (Sinkhorn and Knopp, 1967; Cuturi, 2013).
Broader applicability. As a final remark, the utility of our algorithm extends to graph-based estimators beyond GRFs. For example, it can be used to improve estimates of the PageRank vector, a popular measure of the importance of the nodes of a graph proposed by Page et al. (1998) to rank websites in search engine results. $\sigma$-couplings consistently match or beat the previous best algorithm for coupling walk lengths (Reid et al., 2024c). See Fig. 9 in App. E for the full results.
@@ 4.2 Experiments with $\sigma$-coupled GRFs
We now empirically evaluate $\sigma$-coupled GRFs for variance reduction of graph node kernel estimates. For real-world graphs, we show that lower variance unlocks better approximate inference with scalable graph-based Gaussian processes (GPs), a novel application of GRFs.
@@
Gram matrix approximation. GRFs take the termination probability $p_{\textrm{halt}}$ as a hyperparameter, determining the rate of decay of the geometric distribution over walk length. A smaller value of $p_{\textrm{halt}}$ samples longer walks and gives more accurate kernel estimates, but takes longer to run. The optimal coupling changes depending on $p_{\textrm{halt}}$. We consider the values $p_{\textrm{halt}}\in\{0.1,0.2,0.3,0.4,0.5\}$, finding the optimal permutation $\sigma$ in each case. To find the optimal $\sigma$, we solve the matching problem (App. C.3)
Figure 3: Kernel estimator error vs termination probability. Insets show permutations for $p_{\textrm{halt}}\in\{0.1,0.2,0.3,0.4,0.5\}$.
@@
for a random Erdős-Rényi graph with $N=100$ nodes, taking a permutation order $n=30$ and choosing the $2$-regularised Laplacian kernel as our target. We use the Hungarian algorithm, averaging the cost matrix over every possible node pair $(v_{i},v_{j})\in\mathcal{N}^{2}$. Having computed couplings $\mu^{(\sigma)}$ for each value of $p_{\textrm{halt}}$, we then test the corresponding $\sigma$-coupled GRFs on a variety of real-world graphs. Fig. 3 shows the results for cora ($N=2708$), with the rest left to App. F.1. We plot the relative Frobenius norm error of the Gram matrix approximation $\|\mathbf{K}-\widehat{\mathbf{K}}\|_{\textrm{F}}/\|\mathbf{K}\|_{\textrm{F}}$ with walkers that are i.i.d., antithetic (Reid et al., 2024c) or $\sigma$-coupled. For each $p_{\textrm{halt}}$, $\sigma$-coupled GRFs give equally good or smaller kernel estimator errors. Our OT approach often substantially outperforms antithetic termination: a data-independent, hard-coded algorithm designed specifically to improve GRFs (Reid et al., 2024c). We include visualisations of the optimal permutations for different values of $p_{\textrm{halt}}$ in the inset, verifying that the $\sigma$-coupling adapts to different hyperparamaters.
@@
Novel application: $\sigma$-coupled GRFs for scalable graph-based GPs. We now apply $\sigma$-coupled GRFs to scalable graph-based Gaussian processes (GPs), where improved estimation of the covariance function permits better approximate inference. Scalable GPs are a novel application of GRFs that may be of independent interest (Borovitskiy et al., 2021; Mostowsky et al., 2024).
@@
Consider the task of probabilistic graph interpolation. This aims to predict unknown graph function values, along with principled uncertainty estimates, from an observed set (Pfaff et al., 2020). Take mesh graphs $\mathcal{G}$ where every node $v_{i}\in\mathcal{N}$ has a normal vector $\bm{n}_{i}\in\mathbb{R}^{3}$ (Dawson-Haggerty, 2023). Our task is to predict the $z$-components of a masked set, $\{(\bm{n}_{i})_{z}\}_{i=1}^{N_{\textrm{test}}}$ with $N_{\textrm{test}}=\lfloor 0.05N\rfloor$. To achieve this, we use a graph-based GP with a heat kernel covariance function. We compute a sparse, unbiased approximation of this kernel using GRFs with $\{16,32,64\}$ walkers that are i.i.d., antithetic (Reid et al., 2024c) or $\sigma$-coupled. Details of GP hyperparameter optimisation are given in App. F.2. Fig. 4 shows the results. For mesh graphs of different sizes (the largest as big as $8700$ nodes), we plot the relative Frobenius norm error of the Gram matrix approximation, the test root mean square error (RMSE), and the KL divergence to the true posterior. Our variance reduction method unlocks more accurate predictions and better uncertainty quantification, sometimes by a factor of $>2$.
@@
Figure 4: Graph GP regression. Rows show the kernel estimation error, test set RMSE, and KL-divergence between the true and approximate posterior. Lower is better. $\sigma$-coupled GRFs give better predictions and uncertainty estimates, sometimes by a factor of $>2$. Standard errors are shaded.
@@
Probabilistic interpolation of traffic data. We also train a scalable graph-based GP on a traffic flow dataset of the highways of San Jose, California, curated by Borovitskiy et al. (2021) using data from Chen et al. (2001) and OpenStreetMap. The graph has $1016$ nodes, with speed only known at $325$. We use $250$ randomly-chosen nodes as training data and the remainder as test data. With this sparse, noisy dataset, $\sigma$-coupled GRFs again give substantially better predictions and uncertainty estimates. Fig. 11 in App. F.3 shows the full results.
@@ 5 Discussion and outlook
OT provides a powerful, unifying paradigm for variance reduction with random features. It offers perspectives, proof techniques and numerical algorithms for finding novel RF couplings on continuous and discrete input domains, substantially beating previous algorithms in both settings and with disparate basis functions.
Variance reduction is not all you need. Whilst the presence of variance reduction is unambiguous, downstream benefits tell a more nuanced story. With GRFs for scalable GPs, variance reduction permits much better approximate inference (Sec. 4.2). With RFFs and RLFs, this is not the case. For instance, when approximating attention in Performers (Choromanski et al., 2020), maximising the pointwise kernel estimator variance – the ‘wrong’ OT problem – turns out to improve predictive performance after row normalisation. This shows that, though popular, naive variance reduction is not always the right goal. We believe this to be underappreciated in the literature.
Right framing, wrong cost function. Therefore, we posit that OT provides the right framing for the problem of coupling RFs, but sometimes pointwise kernel variance is the wrong cost function. This choice may not fully capture how the joint distribution over kernel estimates determines downstream performance. Coupling to optimise e.g. the spectral properties of $\widehat{\mathbf{K}}$ (Choromanski et al., 2018; Avron et al., 2017a) or the variance of row-normalised attention scores may prove better. These objectives are rarely considered in the literature. Fortunately, OT provides a suite of theoretical and numerical tools achieve this; one simply modifies the cost function in Eq. 2, optimising a different characteristic of the coupling. We hope this research will spur future work in this exciting direction.
@@ 6 Contributions and acknowledgements
Relative contributions. IR conceptualised the project, proposed the coupling mechanisms in Defs 3.3 and 4.2, proved the major theoretical contributions, ran the GRF experiments and wrote the manuscript. SM designed and ran the RFF and RLF GP experiments (Sec. 3.3), helped shape the project’s direction and made core contributions to the text. KC acted as the senior lead, providing technical guidance and developing the algorithms for the matching problem in App. C.3.1. RET met frequently throughout the project, giving important advice and support. AW provided helpful discussion, supervision and feedback on the manuscript.
Acknowledgements and funding. IR acknowledges support from a Trinity College External Studentship. Part of the work was completed as a student researcher at Google. SM acknowledges funding from the Vice Chancellor’s and the George and Marie Vergottis scholarship of the Cambridge Trust, and the Qualcomm Innovation Fellowship. RET is supported by Google, Amazon, ARM, Improbable and an EPSRC grant EP/T005386/1. AW acknowledges support from a Turing AI fellowship under grant EP/V025279/1 and the Leverhulme Trust via CFI.
@@
7 Ethics and reproducibility
Ethics statement: Our work is foundational with no immediate ethical concerns apparent to us. However, increases in scalability provided by improvements to MC algorithms could exacerbate existing and incipient risks of machine learning, from bad actors or as unintended consequences.
Reproducibility statement: Every effort has been made to ensure the work’s reproducibility. The core algorithms are presented in Defs 3.3 and 4.2, with exhaustive details and discussion in the Appendices. Theoretical results are proved with full assumptions in Apps A and C.3.1, with proof sketches included in the main text for clarity. We will make source code available if accepted. All datasets are available online. We give links to suitable repositories in every instance. Where possible, results are reported with uncertainties to facilitate comparison.